# QT-DoG: Quantization-aware Training for Domain Generalization

### v1

@@ Abstract
Domain Generalization (DG) aims to train models that perform well not only on the training (source) domains but also on novel, unseen target data distributions. A key challenge in DG is preventing overfitting to source domains, which can be mitigated by finding flatter minima in the loss landscape. In this work, we propose Quantization-aware Training for Domain Generalization (QT-DoG) and demonstrate that weight quantization effectively leads to flatter minima in the loss landscape, thereby enhancing domain generalization. Unlike traditional quantization methods focused on model compression, QT-DoG exploits quantization as an implicit regularizer by inducing noise in model weights, guiding the optimization process toward flatter minima that are less sensitive to perturbations and overfitting. We provide both theoretical insights and empirical evidence demonstrating that quantization inherently encourages flatter minima, leading to better generalization across domains. Moreover, with the benefit of reducing the model size through quantization, we demonstrate that an ensemble of multiple quantized models further yields superior accuracy than the state-of-the-art DG approaches with no computational or memory overheads. Our extensive experiments demonstrate that QT-DoG generalizes across various datasets, architectures, and quantization algorithms, and can be combined with other DG methods, establishing its versatility and robustness.  footnotetext: Corresponding author.
CVLab, EPFL
firstname.lastname@epfl.ch
@@ 1 Introduction
In this work, we demonstrate that flatter minima in the loss landscape can be effectively achieved through weight quantization using Quantization-aware Training (QAT), making it an effective approach for DG. By restricting the possible weight values to a lower bit precision, quantization imposes constraints on the weight space, introducing quantization noise into the network parameters. This noise, as discussed in prior works (An, 1996; Murray & Edwards, 1992; Goodfellow et al., 2016; Hochreiter & Schmidhuber, 1994), acts as a form of regularization that naturally encourages the optimization process to converge toward flatter minima. Furthermore, our results show that models trained with quantization not only generalize better across domains but also reduce overfitting to source domains. To the best of our knowledge, this is the first work to explicitly explore the intersection of quantization and domain generalization. Through both theoretical insights and empirical validation, we provide strong evidence that QAT promotes flatter minima, leading to enhanced generalization performance on unseen domains.
@@
  • Quantization for Domain Generalization. We are the first to explore the intersection of quantization and domain generalization, showing that, unlike conventional methods that primarily use quantization for compression, quantization-aware training can significantly improve generalization performance on unseen domains.
  • Flat Minima via QAT: We theoretically and empirically demonstrate that QAT encourages flatter minima, enhancing the generalization of models to unseen domains. Additionally, we show that QAT stabilizes model behavior on OOD data during training.
  • Resource Efficiency: In contrast to traditional DG methods that often increase model size or computational cost, QT-DoG not only improves generalization but also significantly reduces the model size, enabling efficient deployment in real-world applications. EoQ, for instance, requires nearly 6 times less memory than Arpit et al. (2022) and 12 times less training compute compared to Ramé et al. (2023), which trains 60 models for diverse averaging.
  • EoQ We introduce EoQ, a strategy that combines the power of quantization and ensembling for generalization. EoQ achieves state-of-the-art results on the DomainBed benchmark while matching the computational cost and memory footprint as a single full-precision model.
Our code and models will be made publicly available.
@@ 2.1.3 Vision Transformers
Recent studies have increasingly utilized vision transformers for domain generalization (Shu et al., 2023; Sultana et al., 2022). Some approaches enhance vision transformers by integrating knowledge distillation (Hinton et al., 2015) and leveraging text modality from CLIP (Radford et al., 2021) to learn more domain-invariant features (Moayeri et al., 2023; Addepalli et al., 2024; Chen et al., 2024; Huang et al., 2023; Liu et al., 2024).
@@ 2.1.5 Weight Averaging
In contrast to combining or averaging predictions, combining or averaging weights from different training stages or different trained models has also been shown to be robust to OOD data. This approach has been one of the most widely-used techniques in the recent literature (Wortsman et al., 2022b; Matena & Raffel, 2022; Wortsman et al., 2022a; Gupta et al., 2020; Choshen et al., 2022; Wortsman et al., 2021; Maddox et al., 2019; Benton et al., 2021; Cha et al., 2021; Jain et al., 2023; Ramé et al., 2023). For example, SWAD (Cha et al., 2021) uses weight averaging to facilitate finding a flat minimum and shows that it reduces overfitting and generalizes better across distribution shifts. Furthermore, DiWA(Rame et al., 2022b) shows that combining weights from various independently trained model offers more diversity in terms of weight distribution and improves the model’s robustness to OOD data.
Arpit et al. (2022) combines the above mentioned two methodologies by proposing an ensemble of models trained with weight averaging. This is shown to yield superior performance to both weight averaging and ensembling when used independently. However, as mentioned before, it comes at the expense of a large memory footprint and computational cost. In this work, we address this by demonstrating that quantization (model compression) can be used to enhance generalization while reducing the computational cost and memory footprint.
@@ 2.2 Model Quantization
Model quantization is used in deep learning to reduce the memory footprint and computational requirements of deep network. In a conventional neural network, the model parameters and activations are usually stored as high-precision floating-point numbers, typically 32-bit or 64-bit. The process of model quantization entails transforming these parameters into lower bit-width representations, such as 8-bit integers or binary values. Existing techniques fall into two main categories. Post-Training Quantization (PTQ) quantizes a pre-trained network using a small calibration dataset and is thus relatively simple to implement (Nagel et al., 2020; Li et al., 2021; Frantar & Alistarh, 2022; Zhao et al., 2019; Cai et al., 2020; Nagel et al., 2019; Shao et al., 2024; Lin et al., 2024; Chee et al., 2023). Quantization-Aware Training (QAT) retrains the network during the quantization process and thus better preserves the model’s full-precision accuracy. Yang et al. (2023); Esser et al. (2020); Zhou et al. (2017); Bhalgat et al. (2020); Yamamoto (2021); Yao et al. (2020); Shin et al. (2023). In the next section, we provide some background on quantization and on the method we will use in our approach. Our goal in this work is not to introduce a new quantization strategy but rather to demonstrate the impact of quantization on generalization.
@@ 3 Domain Generalization by Quantization
We build our method on the simple ERM approach to showcase the effects of quantization on the training process and on the generalization to unseen data from a different domain.
Despite the simplicity of this approach, we will show in Section 4.3 that it yields a significant accuracy boost on the test data from the unseen target domain. Furthermore, it stabilizes the behavior of the model on OOD data during training, making it similar to that on the in-domain data. In the remainder of this section, we focus on providing some insights on how quantization enhances DG.
@@ 3.2 Quantization Leads to Flat Minima
In the literature (Rame et al., 2022b; Arpit et al., 2022; Krueger et al., 2021; Cha et al., 2021; Rame et al., 2022b; Foret et al., 2021), it has been established that a model’s generalization ability can be increased by finding a flatter minimum during training. This is the principle we exploit in our work, but from the perspective of quantization. In practice, ERM can have several solutions with similar training loss values but different generalization ability. Even when the training and test data are drawn from the same distribution, the standard optimizers, such as SGD and Adam (Kingma & Ba, 2015), often lead to sub-optimal generalization by finding sharp and narrow minima (Keskar et al., 2017; Dziugaite & Roy, 2017; Garipov et al., 2018; Izmailov et al., 2018; Jiang et al., 2020; Foret et al., 2021). This has been shown to be prevented by introducing noise in the model weights during training (An, 1996; Murray & Edwards, 1992; Goodfellow et al., 2016; Hochreiter & Schmidhuber, 1994). Here, we argue that quantization inherently induces such noise and thus helps to find flatter minima.
@@
Eq. 4 shows how the quantization noise $\Delta$ interacts with the curvature $\mathcal{H}$ of the loss function. In regions with large curvature (sharp minima), the Hessian $\mathcal{H}$ has large eigenvalues, and even small perturbations $\Delta$ result in large increases in the loss (Dinh et al., 2017). In contrast, in flat regions (small eigenvalues of $\mathcal{H}$), the loss remains nearly unchanged for small perturbations, encouraging convergence to these flat minima. Therefore, this new loss function with noise induced by quantization encourages model to find a flatter minimum.
In the case of quantization-aware training, the induced noise $\Delta$ is influenced by the quantization bin width or the quantizer step size $s$, and thus ranges between $-\frac{s}{2}$ and $+\frac{s}{2}$. This $s$ is directly dependent on the quantization levels or the bit-width chosen for weight quantization. As the number of bits per weight decreases, the amount of induced noise increases. Hence, the impact of the additional noise can be weighed by choosing an optimal bit-width. As will be shown in Section 4, certain bit-widths thus yield better and flatter minima that enhance generalization. However, if we induce too much noise(very low bit-precision), the weight update rule is dominated by the noise, and the model converge to sub-optimal solutions as shown in the appendix.
@@ 3.3 Empirical Analysis of Quantization-aware Training and Flatness
In this section, we demonstrate that a flatter minimum is reached when incorporating quantization in the ERM process. Our loss flatness analysis shows that QT-DoG can find a flatter minimum in comparison to not only ERM but also SAM (Foret et al., 2021) and SWA (Izmailov et al., 2018).
@@
3.4 Stable Training Process
Figure 3: Model Quantization improves out-of-domain performance as well as training stability. The plots were computed using the TerraInc dataset with domain L100 (left) and L46 (right) as test domain, and the other domains as training/validation data. The network used for these plots was a ResNet-50. For our quantized models, shown in blue in each plot, we quantized the model after 2000 steps. Note that the model accuracy is not only better with quantization but also much more stable for out of distribution data after the quantization step.
3.5 Ensembles of Quantization
For our ensemble creation, we train multiple models and, incorporate quantization into the training process to obtain smaller quantized models. We refer to this as the Ensemble of Quantization (EoQ). As Breiman (1996), we use the bagging method to combine the multiple predictions. Therefore, the class predicted by EoQ for an input $\mathbf{x}$ is given by
4 Experiments
4.1 Datasets and Metrics
PACS (Li et al., 2017) is a 7 object classification challenge encompassing four domains, with a total of 9,991 samples. It serves to validate our method in smaller-scale settings. VLCS (Fang et al., 2013) poses a 5 object classification problem across four domains. With 10,729 samples, VLCS provides a good benchmark for close Out-of-Distribution (OOD), featuring subtle distribution shifts simulating real-life scenarios. OfficeHome (Venkateswara et al., 2017) comprises a total of 15,588 samples. It presents a 65-way classification challenge featuring everyday objects across four domains. TerraIncognita (Beery et al., 2018) addresses a 10 object classification challenge of animals captured in wildlife cameras, with four domains representing different locations. The dataset contains 24,788 samples, illustrating a realistic use-case where generalization is crucial. DomainNet (Peng et al., 2019) provides a 345 object classification problem spanning six domains. With 586,575 samples, it is one of the largest datasets.
4.2 Implementation Details
We use LSQ (Esser et al., 2020) and INQ (Zhou et al., 2017) for model quantization, with the same configuration as existing quantization methods (Esser et al., 2020; Bhalgat et al., 2020; Dong et al., 2019; Yao et al., 2020; Zhou et al., 2017), where all layers are quantized to lower bit precision except the last one. We quantize the models at 8,000 steps for DomainNet and 2,000 steps for the other datasets. Moreover, each channel in a layer has a different scaling factor.
4.3 Results
Table 1: Comparison with domain generalization methods. Performance benchmarking on 5 datasets of the DomainBed benchmark. Highest accuracy is shown in bold, while second best is underlined. do not report confidence interval and ensembles do not have confidence interval because an ensemble uses all the models to make a prediction. Our proposed method is colored in Gray. Average accuracies and standard errors are reported from three trials. For all the reported results, we use the same training-domain validation protocol as (Gulrajani & Lopez-Paz, 2021). $Models$ corresponds to the number of models trained during training and $Size$ corresponds to the relative network size.
@@ 4.3.1 Comparison with DG methods
4.3.2 Bit Precision Analysis
Here, we empirically analyze the effect of different bit-precisions for quantization on the generalization of the model. We perform experiments with four different bit levels and present an analysis in Figure 4 on the PACS (Li et al., 2017) and TerraIncognita (Beery et al., 2018) datasets. We report the test [^1] In-domain results for different bit-precisions are provided in the appendix. domain accuracy averaged across all domains. For both datasets, 7-bit precision was found to be the optimal bit precision to have the best out-of-domain generalization while maintaining in-domain accuracy. Nonetheless, 8 bits and 6 bits also show improvements, albeit smaller than with 7-bit quantization. These results evidence that, even with a 6 times smaller model, quantization still yields better out-of-domain performance without sacrificing the in-domain accuracy.
4.3.3 Combinations with Other Methods
Since QT-DoG requires no modifications to training procedures or model architectures, it is universally applicable and can seamlessly integrate with other DG methods. As shown in Table 3, we integrate QT-DoG with CORAL (Sun et al., 2016) and MixStyle (Zhou et al., 2021). Both CORAL and MixStyle demonstrate improved performance when combined with QT-DoG, reinforcing our theoretical findings that QAT aids in identifying flat minima, thereby enhancing domain generalization.
@@ 4.3.3 Combinations with Other Methods
Algorithm PACS TerraInc C
CORAL 85.5 $\pm$ 0.6 47.1 $\pm$ 0.2 -
CORAL + QT-DoG 86.9 $\pm$ 0.2 50.6 $\pm$ 0.3 4.6x
MixStyle 85.2 $\pm$ 0.3 44.0 $\pm$ 0.4 -
MixStyle + QT-DoG 86.8 $\pm$ 0.3 47.7 $\pm$ 0.2 4.6x

### v2

@@ Abstract
A key challenge in Domain Generalization (DG) is preventing overfitting to source domains, which can be mitigated by finding flatter minima in the loss landscape. In this work, we propose Quantization-aware Training for Domain Generalization (QT-DoG) and demonstrate that weight quantization effectively leads to flatter minima in the loss landscape, thereby enhancing domain generalization. Unlike traditional quantization methods focused on model compression, QT-DoG exploits quantization as an implicit regularizer by inducing noise in model weights, guiding the optimization process toward flatter minima that are less sensitive to perturbations and overfitting. We provide both an analytical perspective and empirical evidence demonstrating that quantization inherently encourages flatter minima, leading to better generalization across domains. Moreover, with the benefit of reducing the model size through quantization, we demonstrate that an ensemble of multiple quantized models further yields superior accuracy than the state-of-the-art DG approaches with no computational or memory overheads. Code is released at: https://saqibjaved1.github.io/QT_DoG/.
Domain Generalization, Quantization, Ensemble, Network Compression, Flat Minima, Regularization
@@ 1 Introduction
In this work, we demonstrate that flatter minima in the loss landscape can be effectively achieved through weight quantization using Quantization-aware Training (QAT), making it an effective approach for DG. By restricting the possible weight values to a lower bit precision, quantization imposes constraints on the weight space, introducing quantization noise into the network parameters. This noise, as discussed in prior works (An, 1996; Murray & Edwards, 1992; Goodfellow et al., 2016; Hochreiter & Schmidhuber, 1994), acts as a form of regularization that naturally encourages the optimization process to converge toward flatter minima. Furthermore, our results show that models trained with quantization not only generalize better across domains but also reduce overfitting to source domains. To the best of our knowledge, this is the first work to explicitly explore the intersection of quantization and domain generalization. Through both analytical reasoning and empirical validation, we provide strong evidence that QAT promotes flatter minima, leading to enhanced generalization performance on unseen domains.
@@
  • We are the first to demonstrate that quantization-aware training, traditionally used for model compression, can serve as an implicit regularizer, with quantization noise enhancing domain generalization.
  • We demonstrate that QAT promotes flatter minima in the loss landscape and provide an analytical perspective behind this effect. Additionally, we show that QAT stabilizes model behavior on OOD data during training.
  • In contrast to traditional DG methods that often increase model size or computational cost, QT-DoG not only improves generalization but also significantly reduces the model size, enabling efficient deployment in real-world applications. EoQ, for instance, requires nearly 6 times less memory than Arpit et al. (2022) and 12 times less training compute compared to Ramé et al. (2023), which trains 60 models for diverse averaging.
@@ 2.1.3 Vision Transformers
Recent studies have increasingly utilized vision transformers for domain generalization (Sultana et al., 2022). Some approaches enhance vision transformers by integrating knowledge distillation (Hinton et al., 2015) and leveraging text modality from CLIP (Radford et al., 2021) to learn more domain-invariant features (Moayeri et al., 2023; Chen et al., 2024; Huang et al., 2023; Liu et al., 2024; Yu et al., 2024; Shu et al., 2023; Addepalli et al., 2024).
@@ 2.1.5 Weight Averaging
Combining or averaging weights from different training stages or models has emerged as a robust approach to improve OOD generalization (Wortsman et al., 2022b; Matena & Raffel, 2022; Wortsman et al., 2022a; Gupta et al., 2020; Choshen et al., 2022; Wortsman et al., 2021; Maddox et al., 2019; Benton et al., 2021; Cha et al., 2021; Jain et al., 2023; Ramé et al., 2023). Techniques like SWAD (Cha et al., 2021) leverage weight averaging to identify flat minima, reducing overfitting and enhancing generalization under distribution shifts. Similarly, DiWA (Rame et al., 2022b) combines weights from independently trained models to improve robustness through increased diversity.
Arpit et al. (2022) integrates ensembling with weight averaging, yielding superior performance compared to either method alone, albeit with significant memory and computational costs. To address these challenges, we demonstrate that quantization can improve generalization while reducing resource demands.
Although flatter minima are not universally indicative of better domain generalization (Andriushchenko et al., 2023), they remain a valuable tool for improving robustness in many scenarios. Moreover, recent findings (Mueller et al., 2023) highlight that selective application of SAM (Foret et al., 2021), such as restricting it to normalization layers, can further refine its effectiveness. The consistent empirical success of SAM underscores its reliability as a method for enhancing domain generalization, despite the nuanced relationship between flatness and performance across different settings.
@@ 2.2 Model Quantization
Model quantization is used in deep learning to reduce the memory footprint and computational requirements of deep network. In a conventional neural network, the model parameters and activations are usually stored as high-precision floating-point numbers, typically 32-bit or 64-bit. The process of model quantization entails transforming these parameters into lower bit-width representations, such as 8-bit integers or binary values. Existing techniques fall into two main categories. Post-Training Quantization (PTQ) quantizes a pre-trained network using a small calibration dataset and is thus relatively simple to implement (Nagel et al., 2020; Li et al., 2021; Frantar et al., 2022; Zhao et al., 2019; Cai et al., 2020; Nagel et al., 2019; Shao et al., 2024; Lin et al., 2024; Chee et al., 2023; Li* et al., 2025; Ramachandran et al., 2024; Shang et al., 2024; Zhang & Shrivastava, 2025). Quantization-Aware Training (QAT) retrains the network during the quantization process and thus better preserves the model’s full-precision accuracy. Yang et al. (2023); Esser et al. (2020); Zhou et al. (2017); Bhalgat et al. (2020); Yamamoto (2021); Yao et al. (2020); Shin et al. (2023). In the next section, we provide some background on quantization and on the method we will use in our approach. Our goal in this work is not to introduce a new quantization strategy but rather to demonstrate the impact of quantization on generalization.
@@ 3 Domain Generalization by Quantization
We build our method on the simple ERM approach to showcase the effects of quantization on the training process and on the generalization to unseen data from a different domain. Despite the simplicity of this approach, we will show in Section 4.1 that it yields a significant accuracy boost on the test data from the unseen target domain. Furthermore, it stabilizes the behavior of the model on OOD data during training, making it similar to that on the in-domain data. In the remainder of this section, we focus on providing some insights on how quantization enhances DG.
@@ 3.1 Quantization
Figure 2: Distribution of Quantization Noise and Weights. We plot the weights(left), quantization noise(middle) and symmetry(right) of random layer in ResNet-50 model. We found KL-divergence to be 0.0009 between quantization noise and uniform distribution with same minimum and maximum value.
3.2 Quantization as uniform noise
Quantization can be modeled as additive noise under certain assumptions (Gray & Neuhoff, 1998). The noise or error introduced by the quantizer is bounded within $\left[-\frac{s}{2},\frac{s}{2}\right]$ where $s$ represents the quantization step size, as described in the previous section. When $s$ is small relative to the dynamic range of the weight distribution, the quantization noise can be well approximated by a uniform distribution (Boncelet, 2009).
This observation is empirically validated in Figure 2, where we analyze the weight distribution of a randomly selected layer and apply 7-bit quantization. We then compute the Kullback-Leibler divergence (Kullback & Leibler, 1951) between the resulting quantization noise and a uniform distribution with the same minimum and maximum values, finding it to be very low. We extend this analysis in Appendix J across multiple layers, we observe that neural network weights generally tend to exhibit smooth and symmetric distributions, reinforcing the conclusion that quantization noise/error generally follows a uniform distribution. Prior works (Zhang et al., 2024; Murray & Edwards, 1992) show that adding noise to network weights can improve generalization, and similarly, quantization-aware training with structured uniform noise also enhances generalization.
@@ 3.3 Quantization Leads to Flat Minima
In the literature (Rame et al., 2022b; Arpit et al., 2022; Krueger et al., 2021; Cha et al., 2021; Rame et al., 2022b; Foret et al., 2021), it has been established that a model’s generalization ability can be increased by finding a flatter minimum during training. This is the principle we exploit in our work, but from the perspective of quantization, and provide an analytical view into how it contributes to achieving flatter minima. In practice, ERM can have several solutions with similar training loss values but different generalization ability. Even when the training and test data are drawn from the same distribution, the standard optimizers, such as SGD and Adam (Kingma & Ba, 2015), often lead to sub-optimal generalization by finding sharp and narrow minima (Keskar et al., 2017; Dziugaite & Roy, 2017; Garipov et al., 2018; Izmailov et al., 2018; Jiang et al., 2020; Foret et al., 2021). This has been shown to be prevented by introducing noise in the model weights during training (An, 1996; Murray & Edwards, 1992; Goodfellow et al., 2016; Hochreiter & Schmidhuber, 1994). Here, we argue that quantization inherently induces such noise and thus helps to find flatter minima. Here, we argue that quantization inherently induces noise, which aids in finding flatter minima. To support this, we use a second-order Taylor series expansion to analyze how quantization-induced perturbations affect the curvature of the loss landscape, leading to flatter minima.
@@
Eq. 4 shows how the quantization noise $\Delta$ interacts with the curvature $\mathcal{H}$ of the loss function. In regions with large curvature (sharp minima), the Hessian $\mathcal{H}$ has large eigenvalues, and even small perturbations $\Delta$ result in large increases in the loss (Dinh et al., 2017). In contrast, in flat regions (small eigenvalues of $\mathcal{H}$), the loss remains nearly unchanged for small perturbations. Quantization noise acts as an implicit regularizer by introducing perturbations $\Delta$ that disrupt the model’s weight updates. In sharper minima, where the Hessian $\mathcal{H}$ eigenvalues are large, small noise significantly increases the loss, causing the model to "escape" these regions and search for flatter, more stable minima. In flatter regions, where the Hessian $\mathcal{H}$ eigenvalues are small, the noise has less impact, helping the model settle into these regions with lower loss. This encourages convergence to solutions that are less sensitive to small changes in the input or model parameters, which is beneficial for out-of-distribution (OOD) generalization.
In the case of quantization-aware training, the induced noise $\Delta$ is influenced by the quantization bin width or the quantizer step size $s$, and thus ranges between $-\frac{s}{2}$ and $+\frac{s}{2}$. This $s$ is directly dependent on the quantization levels or the bit-width chosen for weight quantization. As the number of bits per weight decreases, the amount of induced noise increases. Hence, the impact of the additional noise can be weighed by choosing an optimal bit-width. As will be shown in Section 4, certain bit-widths thus yield better and flatter minima that enhance generalization. However, if we induce too much noise(very low bit-precision), it introduces over-regularization. This excessive noise can overly restrict the search space, preventing the model from reaching a good solution. Instead, the optimization process may focus on minimizing the loss in a way that avoids sharp regions, but sacrifices the ability to find true minimum of the loss function. This is also evident in Table 10 in the appendix.
@@ 3.4 Empirical Analysis of Quantization-aware Training and Flatness
In this section, we demonstrate that a flatter minimum is reached when incorporating quantization in the ERM process. Similar to (Dinh et al., 2017; Cha et al., 2021), we interpret flat minima as "a large connected region in weight space where the error remains approximately constant," as defined by (Hochreiter & Schmidhuber, 1997). Our loss flatness analysis shows that QT-DoG can find a flatter minimum in comparison to not only ERM but also SAM (Foret et al., 2021) and SWA (Izmailov et al., 2018).
@@
Figure 4: Model Quantization improves out-of-domain performance as well as training stability. The plots were computed using the TerraInc dataset with domain L100 (left) and L46 (right) as test domain, and the other domains as training/validation data. The top two plots illustrate in-domain validation accuracy, while the bottom two represent out-of-domain test accuracy. The network used for these plots was a ResNet-50. For our quantized models, shown in blue in each plot, we quantized the model after 2000 steps. Note that the model accuracy is not only better with quantization but also much more stable for out of distribution data after the quantization step.
Table 1: Comparison with domain generalization methods. Performance benchmarking on 5 datasets of the DomainBed benchmark. Highest accuracy is shown in bold, while second best is underlined. do not report confidence interval and ensembles do not have confidence interval because an ensemble uses all the models to make a prediction. Our proposed method is colored in Gray. Average accuracies and standard errors are reported from three trials. For all the reported results, we use the same training-domain validation protocol as (Gulrajani & Lopez-Paz, 2021). $Models$ indicate the number of models trained, and $Size$ represents the relative network size.
@@
3.5 Stable Training Process
3.6 Ensembles of Quantization
For our ensemble creation, we train multiple models independently from initialization, using random seeds to ensure diversity and, incorporate quantization into the training process to obtain smaller quantized models. We refer to this as the Ensemble of Quantization (EoQ). As Breiman (1996), we use the bagging method to combine the multiple predictions. Therefore, the class predicted by EoQ for an input $\mathbf{x}$ is given by
4 Experiments
We evaluate our approach on diverse datasets from Domainbed and WILDS (Koh et al., 2021) Benchmark. All implementation, datasets, metric details and various ablation studies are provided in the Appendix.
4.1 Results
@@ 4.1.1 Comparison with DG methods
Algorithm PACS TerraInc C
CORAL 85.5 $\pm$ 0.6 47.1 $\pm$ 0.2 -
CORAL + QT-DoG 86.9 $\pm$ 0.2 50.6 $\pm$ 0.3 4.6x
MixStyle 85.2 $\pm$ 0.3 44.0 $\pm$ 0.4 -
MixStyle + QT-DoG 86.8 $\pm$ 0.3 47.7 $\pm$ 0.2 4.6x
Table 3: Comparison between ERM and QT-DoG on the Amazon and Camelyon datasets. We report the in-domain and out-of-domain accuracy with respective metrics as shown. C represents the compression factor of the model.
Dataset Method In-dist Out-dist C Metric
Amazon ERM 71.9 $\pm$ 0.1 53.8 $\pm$ 0.8 - 10th percentile acc
Amazon QT-DoG 79.2 $\pm$ 0.5 55.9 $\pm$ 0.6 4.6x 10th percentile acc
Camelyon ERM 93.2 $\pm$ 5.2 70.3 $\pm$ 6.4 - Average acc
Camelyon QT-DoG 96.4 $\pm$ 2.1 78.4 $\pm$ 2.2 4.6x Average acc
@@
4.1.2 Combinations with Other Methods
Since QT-DoG requires no modifications to training procedures or model architectures, it is universally applicable and can seamlessly integrate with other DG methods. As shown in Table 2, we integrate QT-DoG with CORAL (Sun et al., 2016) and MixStyle (Zhou et al., 2021). Both CORAL and MixStyle demonstrate improved performance when combined with QT-DoG, reinforcing our findings that QAT aids in identifying flat minima, thereby enhancing DG.